### v1

# Co-VisiON: Co-Visibility ReasONing on Sparse Image Sets of Indoor Scenes

Abstract
Humans exhibit a remarkable ability to recognize co-visibility—the overlapping regions visible in multiple images—even when these images are sparsely distributed across a complex scene. This capability is foundational in 3D vision and robotic perception. Despite significant progress in vision learning, it remains unclear whether current vision models have reached human-level proficiency in co-visibility analysis. In this work, we introduce the Co-Visibility reasONing (Co-VisiON) benchmark, designed to directly evaluate co-visibility reasoning on sparse image sets across over 1000 indoor scenarios. Our experiments reveal that while co-visibility is typically treated as a low-level feature matching task, it poses a significant challenge for existing vision models under sparse conditions. Notably, a proprietary vision-language model outperforms all purely vision-based approaches, with all models lagging substantially behind human performance. This gap underscores the need for more than basic pairwise vision processing—it calls for a comprehensive spatial understanding through high-level reasoning across multiple views. Inspired by human visual cognition, we propose a novel multi-view baseline, Covis, which achieves top performance among pure vision models and narrows the gap to the proprietary VLM. We hope our benchmark and findings will spur further advancements in developing vision models capable of robust, high-level reasoning in challenging, sparse environments. Our dataset and source code can be found at https://ai4ce.github.io/CoVISION.
1 Introduction
Humans are remarkably adept at constructing cognitive topology maps from multi-view observations, enabling them to accurately infer a scene’s layout. For example, when browsing websites for real estate listings or hotel bookings, one is often presented with only a few spatially sparse images. Despite the limited visual data, humans can intuitively grasp the spatial context and deduce the relationships between images, effectively constructing a mental graph where images serve as nodes and co-visible regions form the connecting edges.
Figure 1: Co-VisiON overview. If two images share non-zero co-visible surface areas, they should be connected in the binary co-visibility graph. Note that the bird’s-eye view is only for visualization and is not actually used in Co-VisiON.
We refer to this ability as co-visibility reasoning because the key step in this process is to look for overlapping co-visible surface areas between the images. Such areas are qualitatively illustrated in Fig. 1 with colored masks on the images on the left along with the colored field-of-views on the BEV map on the right. Co-visibility also plays a crucial role in 3D computer vision and robot perception, underpinning applications such as image matching, scene reconstruction, visual localization, and mapping.
Modern vision models have achieved remarkable progress in these tasks and have been proficient in matching co-visible images when they have dense overlaps. However, we are particularly intrigued by humans’ robust co-visibility reasoning in sparsely distributed images, where shared visible information is limited or highly distorted. Can current vision models compete with human proficiency in co-visibility reasoning under sparse conditions?
Figure 2: Complex Co-VisiON Examples. This figure showcases Co-VisiON in two distinct scenarios, each featuring a selected pair, or pairs, of co-visible images that reflect explicit reasoning. On the left, weak feature correspondence and sparse angles are challenging to reason. On the right, the connection between images A and B is not obvious until the introduction of image C. These examples demonstrate the challenge in co-visibility reasoning beyond feature matching between image pairs and call for reasoning on multi-views simultaneously. The pairwise processing falls short when facing diverse viewpoints with significant angular and spatial disparities. Additional views can offer contextual information from the scene’s periphery, making it easier to overcome these challenges.
In this work, we propose the Co-VisiON benchmark, which evaluates models on Co-Visibility reasONing from sparsely captured images across over 1,000 indoor scenarios. We benchmark a diverse set of relevant methods, including feature matching [33, 18, 44, 53], contrastive learning [10], visual place recognition [4], 3D reconstruction [58, 55], supervised graph edge prediction [15, 21, 50] and large vision-language models [39, 64]. However, our results indicate that these methods fail to capture the broader context and spatial relationships within a scene.
As shown in Fig. 2, images with minimal overlap and sparse views present a significant challenge. This highlights the need for more holistic scene understanding beyond local feature matching. Beyond its challenges, co-visibility reasoning is a fundamental capability for many 3D vision applications. Having a reliable co-visibility graph can significantly improve the efficiency of structure-from-motion (SfM) [65] and simultaneous localization and mapping (SLAM) [37] under sparse observations. Similarly, it can benefit robotics tasks such as place recognition, navigation, and pose estimation, where reasoning about spatial relationships from limited viewpoints is crucial [24].
Besides the aforementioned baselines, we propose our baseline, Covis, a supervised binary classification and mask segmentation method with either pairwise or multiple views as input.
We notice that the multi-view approach, capturing more scene information, outperforms pairwise methods, provided that the additional views did not contain irrelevant information (i.e., non-co-visible features). Inspired by the confidence-level in DUSt3R [58], we propose a learnable mask supervised by co-visible regions. Applying this mask to the features filters out irrelevant information and improves co-visibility reasoning by focusing on the most relevant features.
In summary, our contributions are:
  • We propose co-visibility reasoning as a novel vision task to assess an intelligent agent’s spatial reasoning ability given sparse observations, and generate a sparse indoor image dataset with dense co-visibility annotations from photorealistic indoor simulation environments.
  • We benchmark a diverse range of baselines (including human) and demonstrate the challenge and potential of advancing co-visibility reasoning research.
  • We demonstrate that in supervised co-visibility reasoning, the multi-view method outperforms two-view ones when assisted with co-visible mask segmentation, indicating the importance of multi-view analysis in this task.
2 Related Work
2.1 Related tasks
Visual Reasoning. Visual reasoning tasks primarily focus on object-level relationships within a single image, such as visual question answering (e.g., GQA [23], CLEVR [26]), and scene graph generation [63, 25]. ISVQA [6] extends this to multi-image reasoning in a household setting, but it mainly targets object relationships, while Co-VisiON emphasizes image relationships grounded in the environment.
Visual Place Recognition. Visual place recognition (VPR) involves identifying previously seen locations and is widely studied in robotics and autonomous vehicles [34, 47, 12, 9].
The prominent VPR method, NetVLAD [4], treats VPR as an image retrieval problem, representing image features as a bag of visual words. The key distinction between VPR and co-visibility reasoning is in spatial requirements: VPR requires geographically close images, while Co-VisiON links images by overlapping surfaces regardless of distance or camera pose. Our experiments show that VPR-based methods like NetVLAD are insufficient for Co-VisiON.
Visual Localization and 3D Reconstruction. Visual SLAM simultaneously performs image-based localization and 3D mapping from (at least) RGB cameras [5, 16, 27, 36, 35, 51], whereas Structure-from-Motion [65, 46, 40] fully reconstructs 3D scenes upto a scale and estimates camera parameters in an offline fashion. In contrast, Co-VisiON builds a co-visibility graph by linking images of a scene based on shared surface visibility, without requiring full scene reconstruction or camera pose estimation.
Visual Correspondence. Estimating visual correspondence between images has wide applications in localization [37], image matching [44], and scene understanding [19]. While traditional methods focus on establishing geometric correspondence such as points [44, 41, 24], lines [65], and planes [2, 29], recent approaches focus on semantic correspondence, such as object relationships [19, 20, 17]. Notably, Rau et al. [43] proposes an image embedding method for estimating visual surface overlap—a task related to ours but limited to image pairs. However, this method performs poorly on the Co-VisiON benchmark (only 20% graph IoU). Overall, such visual correspondence tasks operate on image pairs and rely on geometric metrics, making it fundamentally different from Co-VisiON that reasons over sparse image sets with significant pose variation and is evaluated using graph IoU.
Related datasets. Benchmarking Co-VisiON requires sparse image sets with ground truth co-visibility graphs. However, many existing real-world datasets [13, 66, 30, 43, 49], including those used in [43], either do not provide such graphs or make it difficult to compute co-visible regions accurately, as those derived from point clouds or depth data are often too coarse and noisy. To obtain high-quality co-visibility ground truth, a natural solution is to leverage virtual 3D scenes with textured meshes, which can be rendered from arbitrary viewpoints with pixel-level co-visibility via OpenGL. Using photorealistic simulation environments like Gibson [62] and HM3D [42], the Co-VisiON dataset provides a sparse yet well-distributed set of viewpoints with full scene coverage and accurate co-visible correspondences, making it well-suited for advancing co-visibility reasoning.
2.2 Related methods
Feature matching. Feature matching approaches range from using handcrafted descriptors like SIFT [33] with RANSAC [18] to more recent deep-learning methods like [53, 44]. SuperGlue [44] employs a graph neural network to establish correspondences by solving an optimal transport problem on deep-learning based descriptors like SuperPoint [14]. In this work, we show such naive pairwise matching methods are not appropriate to solve Co-VisiON.
Image Representation Learning. Image representations can be learned through contrastive objectives, pairwise supervision, or spatially guided pretraining. Contrastive methods [38, 10, 22, 61, 56] bring similar images closer while pushing dissimilar ones apart. NetVLAD [4] extends this by aggregating local features into global descriptors. Pairwise classification methods predict whether two images depict the same place, typically using different backbones [15, 50, 21, 60]. While effective in simple settings, these approaches operate on image pairs and miss broader context. Alligat0R [32], as an example of spatially guided pretraining, and Covis both use co-visibility signals, but for different goals: Alligat0R uses them for pose regression, while we use them to mask irrelevant features in multi-view co-visibility prediction. The approaches above were not designed for scene-level reasoning, and our findings suggest they struggle to capture the spatial understanding.
Reasoning with vision and language models. Notably, both image matching and representation learning focus on correspondences between images from similar viewpoints or the same image with augmentations. In contrast, co-visibility reasoning requires matching images from significantly different viewpoints, demanding spatial understanding beyond feature comparison. Since VLMs [7, 39] demonstrate reasoning abilities [28, 8], we explore their potential by prompting GPT-4o [39] and Qwen-72B [1] for co-visibility reasoning between multiple images.
Multi-view learning. Multi-view learning [67] explores how information from multiple images can be integrated for tasks like object-level classification [48, 52] and 3D reconstruction [55]. Notably, MV-DUSt3R [55] fuses multi-view images to generate accurate 3D geometry with an emphasis on metric-level detail. In contrast, few works have explored multi-view learning for high-level scene understanding, which involves inferring the topological structure of a scene from sparse observations. Co-VisiON tackles this underexplored setting by jointly reasoning over multiple images to construct a co-visibility graph.
3 Co-visibility Benchmark
3.1 Task Definition
Given a sparse set of images =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$ from a scene representing different views of a house, our objective is to construct a co-visibility graph $\mathcal{G}$. In the graph, images serve as nodes and they are connected by an edge only if they have co-visibility. Specifically, for each pair of images $(I_{i},I_{j})$, an algorithm outputs $d_{ij}$ that quantifies the degree of co-visibility between the two images.
Next, we employ thresholding to determine if $d_{ij}$ surpasses a predefined threshold $\tau$. If $d_{ij}>\tau$, we connect $I_{i}$ and $I_{j}$ with an edge in the binary graph $\mathcal{G}$. Covering all possible pairs of images, we construct an adjacency matrix $\mathbf{A}$, where each entry $a_{ij}$ is binarized from $d_{ij}$ by the threshold:
$a_{ij}=1\iff d_{ij}>\tau,$ (1)
indicating co-visibility between $I_{i}$ and $I_{j}$.
The resulting adjacency matrix $\mathbf{A}$ is the final representation of the co-visibility graph $\mathcal{G}$. Therefore, by evaluating the matrix $\mathbf{A}$, a model’s ability to reason co-visibility from the sparse image set can be quantitatively measured.
3.2 Evaluation metrics
Following the notation introduced in Sec. 3.1, we denote the predicted co-visibility graph and adjacency matrix by a model as $\hat{\mathcal{G}}$ and $\hat{\mathbf{A}}$ respectively. To evaluate the $\hat{\mathbf{A}}$ against the ground truth $\mathbf{A}$ of graph $\mathcal{G}$, we use two evaluation metrics; Intersection over Union (IoU) and Area Under Curve (AUC), as following:
Graph Intersection over Union (Graph IoU): Graph Intersection over Union (Graph IoU) measures the similarity between $\hat{\mathcal{G}}$ and $\mathcal{G}$ by comparing the overlap of their edges [68]. It is defined as the ratio of common edges to the total number of unique edges in both graphs:
Graph $\text{Graph IoU}=\frac{|\hat{\mathbf{A}}\cap\mathbf{A}|}{|\hat{\mathbf{A}}\cup\mathbf{A}|}.$ (2)
Area Under Curve (AUC): Since graph IoU scores can be sensitive to the chosen thresholds, we further report the Area Under Curve (AUC) to ensure fairness and robustness in task evaluation. AUC is obtained by plotting the graph IoU for different thresholds sampled evenly from the range $[0,1]$. The motivation for using AUC is that a good co-visibility reasoning method should not just achieve good IoU for a specific threshold, but also achieve consistently high IoU scores on a wide range of thresholds. A higher AUC indicates that the model is consistently better and more robust. This will particularly benefit real-world applications where threshold optimization may not be feasible. The evaluation metrics are detailed in the supplementary.
3.3 Dataset Generation
To construct a structured multi-view indoor dataset with well-defined co-visibility relationships, we employ a progressive candidate selection strategy for camera placement. Each image is selected to contribute to the overall scene representation. The selection is based on the highest intersection over union (IoU) with previously chosen images, ensuring the IoU is neither too high ($\geq$30%, reducing sparsity) nor too low ($\leq$5%, leading to a disconnected graph). This iterative process continues until a predefined coverage ratio is achieved. The resulting dataset consists of a sparsely distributed image set, with a co-visibility graph encoding pairwise relationships and corresponding co-visible region information (detailed in Sec. 3.1). We use Habitat-sim [54, 45] to collect sparse yet representative data from Gibson [62] and HM3D [42], with scene-level splits to prevent data leakage.  Tab. 1 presents dataset statistics, and Fig. I illustrates the full dataset pipeline, with additional details provided in the supplementary material.
Gibson. Gibson provides a collection of 3D models of real-world indoor environments, encompassing various types such as homes and offices. We select scenes based on the availability of .glb files and necessary metadata. Depending on the scene’s scale, the number of images ranges from a dozen to over one hundred. On average, approximately 14 images are captured per scenario for each of the 85 distinct scenes, resulting in a comprehensive dataset used for co-visibility graph annotation. The Gibson dataset is split into 80% for training and 20% for testing.
HM3D. We generate the second half of the Co-VisiON dataset from HM3D. A key difference between Gibson and HM3D is the availability of floor height data required for scene construction. While Gibson encodes floor heights in its metadata, HM3D lacks this information. To address this, we use navigable regions to distinguish floor heights and define the scene’s vertical structure. The HM3D dataset is divided into 90% for training and 10% for testing.
Table 1: Statistics of Co-VisiON Dataset. Scenarios is defined as structured collections of images with associated camera poses within a given scene.
Gibson HM3D
1-floor scenes 50 257
2-floor scenes 27 309
3-floor scenes 7 161
$\geq$4-floor scenes 1 28
Scenarios / scene 5 1
Total scenes 85 755
Total scenarios 425 755
Total images 5,954 16,245
Total image pairs 33,849 210,008
Human Annotations. We collect human annotation as part of the Co-Vision dataset, serving as a baseline in Co-VisiON benchmark. A binary graph is generated by several experienced annotators for each scenario in Gibson. It is then compared to the ground truth to compute the graph IoU. This process represents the level of human reasoning capability for Co-VisiON.
4 Experiments on Co-VisiON
Figure 3: Multi-view Covis pipeline. Multi-view Covis is built on the MV-DUSt3R [55] encoder-decoder backbone, following the same attention pattern across views. Visual tokens for the reference, positive, and negative views are shown in blue, green, and orange respectively. Gray and bold arrows indicate cross-view attention in the decoder. The reference (anchor) uses a separate prediction head, while the positive and negative views share the classification (CLS) and mask heads. Decoder outputs are projected to pixel-wise features $F\in\mathbb{R}^{H\times W\times C}$, then filtered by a mask $M$ to obtain $F^{\prime}$.
4.1 Naive Baselines
We first benchmark several conventional approaches for Co-VisiON, where inference involves exhaustive pairwise comparisons between all node pairs to generate a topology map. This includes binary classification, supervised contrastive learning and feature matching.
Pairwise supervised classification. We test three classic image classification backbones, including ViT [15], VGG [50], and ResNet [21]. They extract representations for a pair of images and append a classification head to predict binary co-visibility probabilities. The models are fine-tuned end-to-end. For inference, we threshold the pairwise prediction probabilities to generate a binary adjacency matrix for each scenario. The area under the curve (AUC) is calculated at various thresholds ranging from 0 to 1, as detailed in Sec. 3.2.
Contrastive learning approaches. We test two contrastive approaches. The first adopts a ResNet backbone with the InfoNCE loss [38] following the architecture of SimCLR [10]. Unlike SimCLR, we conduct supervised contrastive learning. Pairs of images connected in the co-visibility graph are treated as positive pairs, and others as negative pairs. The model is fine-tuned end-to-end from the pretrained SimCLR. The second approach utilizes NetVLAD [4], a widely used method in visual place recognition. It uses a ResNet backbone followed by a VLAD aggregation layer to extract global image descriptors. For both approaches, the pairwise feature distance, normalized by the maximum feature distance (ranging from 0 to 1), is used during inference to predict the edge probability for the co-visibility graph.
Image feature matching. SIFT+RANSAC [18] and SuperGlue [44] are used as baseline methods for image feature matching due to their classic and widespread use. SuperGlue processes image pairs to detect key points and compute matches via a graph neural network, while SIFT+RANSAC employs traditional feature matching and outlier filtering. The percentage of matches compared to the maximum possible matches is treated as the edge probability for the co-visibility graph.
Discussion. We present the experiment results in  Tab. 2. Compared with only reporting the best IoU, we also emphasize the importance of IoU invariance across various thresholds as a critical performance metric, particularly in real-world applications where real-time threshold optimization may not be feasible. A robust baseline should demonstrate consistent performance across a wide range of thresholds, ensuring acceptable results regardless of the threshold chosen. In Fig. 4, we plot the AUC curves of all baselines. Vision Transformer (ViT) stands out for its stable IoUs across diverse thresholds among all the conventional baselines. Contrastive learning shows a similar trend. Comparing the baselines by their categories, supervised learning, particularly with ViT, consistently outperforms feature-matching methods in both datasets shown in Tab. 2. Contrastive learning methods, although trailing behind supervised methods, still outperform feature-matching techniques by effectively managing positive and negative pairs in feature space. This showcases the great potential of contrastive learning in addressing co-visibility issues. All conventional baselines are substantially worse than the human evaluation. Pairwise image matching and classification baselines are not sufficient to solve Co-VisiON.
4.2 Improved Baselines
Going beyond conventional baselines, we explore whether recent vision models can perform better on Co-VisiON, including state-of-the-art 3D stereo reconstruction methods that can inherently derive co-visibility, and large Vision and Language Models (VLMs) that excel in visual understanding. More importantly, we propose and test a new baseline–Covis, leveraging multi-view masked classification.
VLM for Co-VisiON. We leverage external API calls to harness the reasoning capabilities of state-of-the-art large vision and language models (VLMs). This method encodes images into base64 format and sends them to GPT-4o [39] and Qwen2.5-VL 72b [64] models. Then they use their pre-trained capabilities to compare the images, compute pairwise scores, and generate a topology map. We further include results from recent Spatial VLM methods like SpatialRGPT [11] and Gemini-2.0-Flash [57] in the appendix.
3D Reconstruction for Co-VisiON. We use the DUSt3R and MV-DUSt3R+ frameworks to generate global point clouds from input photos. The presence of binary topology edges is determined by the overlap of these point clouds. This approach is counterintuitive, as 3D reconstruction typically relies on topology map node relationships. Instead, we use 3D reconstruction to generate the topology map to highlight its limitations, including domain gap issues and frequent misjudgments of negative samples shown in Tab. 2.
4.3 Covis
By analyzing human evaluation patterns, we design Covis, a multi-view encoder-decoder pipeline for co-visibility prediction, shown in Fig. 3. To validate the advantages of multi-view reasoning, we perform an ablation study comparing Covis with its pairwise version. This helps assess the benefits of leveraging multiple views.
Architecture. Both pairwise and multi-view Covis use CroCov2 [60] as the backbone with a BCE loss to predict the relationship between images. In the pairwise setting, a reference image is paired with a single image from another viewpoint, and the BCE loss predicts whether the two images are co-visible. In contrast, the multi-view Covis is built on the MV-DUSt3R [55] encoder-decoder backbone that facilitates the exchange of information between multiple images from diverse viewpoints, allowing for richer spatial reasoning. During training, positive and negative pairs are constructed based on the ground truth co-visibility graph, and during inference, a mixture of positive and negative pairs from the same scene is used to assist pairwise reasoning. The multi-view approach captures broader context by processing multiple images simultaneously.
Mask prediction. We found that non-covisible information contaminates the learned representation of the reference image, hindering effective learning from positive views (see Appendix for details). To address this, we introduce a learnable mask $M\in\mathbb{R}^{H\times W}$, jointly optimized as an auxiliary objective and used to filter out irrelevant features.
$\mathbf{M}=\operatorname{conv}(\mathbf{F}),$ (3)
where $F\in\mathbb{R}^{H\times W\times C}$ denotes the pixel-wise feature map obtained by projecting the decoder output. After obtaining $\mathbf{M}$, we apply it to $\mathbf{F}$ via a dot product at each pixel, preserving relevant information from the positives:
$\mathbf{F}^{\prime}=\mathbf{F}\odot\mathbf{M},$ (4)
where $\odot$ denotes pixel-wise multiplication. Both the topology generation and mask prediction tasks use BCE loss but serve distinct purposes: the former predicts spatial relationships based on flattened 1D global features for image pairs, while the latter supervises pixel-wise co-visible regions, providing stronger localized signals.
(a) IoUs for Gibson dataset
(b) IoUs for HM3D dataset
Figure 4: Comparison of IoU scores across various thresholds for the top-performing vision models on the Gibson and HM3D datasets.
4.4 Discussion
Table 2: IoU* (best IoU) and AUC results on the Gibson and HM3D datasets. We compare baseline and manual performance on the test sets. Bold values indicate the best performance in each category. The top three performances are color-coded: red for the best, teal for second, and blue for third. For 3D reconstruction, VLM categories and human annotation, AUC values are equivalent to IoU since they are not affected by thresholding. Binary CLS. stands for binary classification. Human annotation performance represents the upper bound.
Category Method Backbone Pairwise(✗) / Multiview(✓) Gibson HM3D
IoU* ($\uparrow$) AUC ($\uparrow$) IoU* ($\uparrow$) AUC ($\uparrow$)
Feature Matching SuperGlue [44] 0.47 0.11 0.38 0.10
SIFT [33] + RANSAC [18] 0.35 0.05 0.34 0.05
Contrastive - ResNet18 [21] 0.35 0.33 0.35 0.29
NetVLAD [4] ResNet18 [21] 0.38 0.32 0.42 0.39
3D Reconstruction DUSt3R [58] CroCo v2 [60] 0.54 0.54 0.40 0.40
MV-DUSt3R+ [55] CroCo v2 [60] 0.56 0.56 0.45 0.45
Binary CLS. - ViT [15] 0.47 0.43 0.48 0.46
- VGG [50] 0.35 0.30 0.34 0.21
- ResNet18 [21] 0.38 0.36 0.48 0.46
- CroCo v2 [60] 0.50 0.45 0.43 0.38
- CroCo v2 [60] 0.48 0.41 0.41 0.37
Masked Binary CLS. Covis (Ours) CroCo v2 [60] 0.56 0.54 0.53 0.51
Covis (Ours) CroCo v2 [60] 0.59 0.57 0.57 0.56
VLM GPT-4o [39] 0.63 0.63 0.59 0.59
GPT-4o [39] 0.58 0.58 0.54 0.54
Qwen2.5-VL 72b [64] 0.41 0.41 0.39 0.39
Human Annotation 0.72 0.72 - -
Pairwise vs multi-view performance. Across both 3D reconstruction and Covis with mask prediction, we observe that multi-view approaches consistently outperform pairwise approaches by about 3% in IoU, as shown in  Tab. 2. Multi-view inputs provide richer spatial context, leading to more accurate 3D reconstruction and better feature learning. In contrast, pairwise Covis is limited by the information in a single image pair, resulting in reduced performance.
The importance of masking. As discussed, multi-view methods typically outperform pairwise methods. However, our experiments with the non-masking version of Covis reveal that their performance degrades sharply as the proportion of negatives increases, particularly in multi-view settings. This is due to feature contamination, where irrelevant features from positives and negatives interfere during information exchange. Inspired by the confidence mechanism in DUSt3R [58], we introduce a learnable mask to filter out irrelevant information. Our results show that applying this mask improves performance by approximately $10\%$, with especially pronounced gains in multi-view scenarios.
VLM performance. We test both proprietary and open-source models without fine-tuning. GPT-4o achieves 63% IoU, which is currently the closest to the human reasoning performance of 72%, top-performed in  Tab. 2. This suggests that reasoning-capable models can better identify co-visibility, especially under the sparse conditions of the Co-VisiON benchmark. In contrast, the open-source Qwen2.5-VL 72b performs less favorably, reaching only 41%, even lower than many visual baselines. This highlights the instability of VLMs on unseen tasks, despite their promising potential for vision understanding.
5 Applications of Co-VisiON
5.1 DUSt3R for scene reconstruction
Co-VisiON guided 3D Reconstruction. Topology maps guide 3D reconstruction by selecting pertinent frames within a specific area. For instance, when reconstructing a kitchen, only frames depicting the kitchen are considered relevant. In contrast, DUSt3R [58] exhaustively uses all available frames regardless of spatial relevance, resulting in inefficiencies in both memory and computation. To improve the accuracy and efficiency of 3D reconstruction, we propose a sparse co-visibility graph and compare it against commonly used topologies such as random, star, complete, and the ground truth. Definitions of these graph structures are provided in the supplementary material. Specifically, we treat each frame as a reference and jointly optimize it only with the frames directly connected in the topology graph, thereby constraining the matching and reconstruction.
Table 3: Comparison of different graph structures within the DUSt3R framework. We evaluate the Average Pose Error (APE), along with time complexity (in seconds) and memory cost (in GB) per scene. Results are averaged over the Gibson test set. The Random graph fails in the 3D reconstruction test, making APE not applicable.
Method Random Star Covision Complete GT
APE (meters)($\downarrow$) Failed 1.791 1.701 1.601 1.548
Time Complexity($\downarrow$) - 23 21 32 25
Memory Cost($\downarrow$) - 7.3 6.7 10.2 7.9
Quantitative Analysis of Different Graphs. We calculate the average pose errors (APE) by matching estimated poses with their closest ground truth poses to evaluate pose estimation and reconstruction quality, as shown in Tab. 3. As expected, using complete and ground truth (GT) graphs yields lower APE, demonstrating more accurate 3D reconstruction. However, obtaining a ground truth graph and using the complete graph incurs high annotation and computational costs. Moreover, random and star graphs fail to produce a reliable 3D map, as shown in Fig. III. Therefore, we find the co-visibility graph to be one of the most suitable options for DUSt3R, balancing memory costs with a reliable 3D map, as shown in Fig. 5. Detailed reconstruction visualizations are available in the supplementary.
(a) Star - Scene 1
(b) Covision - Scene 1
(c) GT - Scene 1
(d) Star - Scene 2
(e) Covision - Scene 2
(f) GT - Scene 2
Figure 5: 3D reconstruction results by DUSt3R using co-visibility and GT graphs, showing similar performance.Regions highlighted by red boxes indicate reconstruction failures.
5.2 Automated Training Set Labeling
The co-visibility graph provides a topology for automatically labeling training data, such as selecting triplets in VPR, forming batches in 3D reconstruction, and pairing images for view completion. We illustrate this with CroCo [59], a model that completes masked regions of a target image using a reference from a different viewpoint.
To eliminate the need for manual annotation, we automate the process by leveraging graphs, where images are represented as nodes, and the edges encode co-visibility relationships based on 3D structure or estimated visibility. This approach allows for the automatic generation of training pairs by treating each image as a query and its connected neighbors as reference candidates.
We evaluate three strategies for generating CroCo training pairs: (1) a high co-visibility graph with over 50% 3D surface overlap; (2) a sparse co-visibility graph from our method; and (3) a random graph. For each, masked regions are restricted to the co-visible area between query and reference, except in the random graph where masks are applied arbitrarily. This allows us to assess whether our sparse co-visibility graph is a more annotation-efficient alternative while maintaining performance.
Table 4: PSNR for cross-view completion with different graphs.
Method Random Co-Vis High Co-Vis
PSNR ($\uparrow$) 12.46 16.34 16.32
Quantitative Analysis on different graphs. We conduct completion quality tests as shown in Fig. 6. Our results show that the sparse co-visibility graph achieves performance comparable to the high co-visibility graph as shown in Tab. 4, while being significantly easier to construct. This is because the sparse graph only requires selecting image pairs that meet minimal co-visibility criteria, whereas the high co-visibility graph demands stricter thresholds and more precise filtering, making it more computationally intensive and harder to scale. Moreover, it generalizes better to diverse scenes and supports more scalable training. These advantages make it a practical and efficient solution for automated training set labeling.
Reference
Prediction
Target
Figure 6: Visualizations of the reconstructed co-visibility regions during inference. The top row shows model inference using the co-visibility graph for training, while the bottom row uses the high co-visibility graph. The three columns represent the reference image, predicted image, and ground truth target image. In the predicted images, the masked regions are visible as network input, and the brightened areas indicate the reconstructed output. Additionally, the random graph fails in cross-view completion.
6 Conclusion
Limitation. The current co-visibility reasoning task presents several notable limitations. For the dataset, the scenes are restricted to the Gibson and HM3D environment, and the human annotations are limited. With the proposed data collection pipeline and website, our future work will continue to enlarge and diversify the dataset scenes with other sources, including the real world.
Summary. This paper proposes the Co-Visibility reasONing (Co-VisiON) task, where models are presented with sparse image sets taken in an indoor environment and are tasked with reasoning their co-visibility relationships by determining view connectivity and predicting a binary co-visibility graph. We curate a dataset designed for this task, containing both human and automated annotations, and conduct comprehensive experiments to show how the Co-VisiON task challenges dominant image classification and matching baselines. Additionally, we introduce Covis, which is able to leverage multi-view reasoning to effectively capture complex spatial relationships between images, outperforming traditional pairwise approaches. We hope that the Co-VisiON task can facilitate future research in embodied scene understanding.
Acknowledgement. We thank Kevin Lee, Kingston Yip, and Mandy Mao for their early investigation on the dataset and baselines, and Leon Zhang and Ammaar Firozi for their contributions to the dataset. The work was supported in part through NSF grants 2238968, 2322242, 2121391, and 2036870, and the NYU IT High Performance Computing resources, services, and staff expertise.
Appendix
Figure I: Dataset generation: Firstly, candidates are sampled from the possible locations shown as red border of the top-down map in candidate selection. Then the sampled candidates, along with a scene coverage map and the current co-visibility graph, are used to calculate scores for each candidate. The scores are computed with the best saved candidates from previous iterations. The top-scoring candidate among the currently sampled is chosen to update scene coverage and the co-visibility graph. This process is repeated until a scene coverage threshold is reached, at which point the best candidates’ observations with a co-visibility graph are saved to create the dataset.
This supplementary provides additional details and results that could not fit in the main paper. Specifically, we include: (1) dataset and annotation generation (Gibson and HM3D), (2) top-down visualization of co-visibility graph, (3) evaluation metric details, (4) graph definitions used in DUSt3R and CroCo experiments, (5) ablation studies on Covis, (6) additional models and failure mode analysis, and (7) Sim2Real downstream task details. All experiments and dataset generation were conducted on an HPC cluster with A100 and V100 GPUs, using a single GPU per run.
Figure II: We are able to generate different scenarios for a single scene. This is an example of 5 scenarios generated for Scene Goff
A Dataset details
We detail the dataset construction and annotation processes for the Gibson and HM3D environments, which form the basis of our co-visibility analysis. For each dataset, we define scenarios as structured collections of images with associated camera poses, designed to ensure high coverage and annotation quality. Both datasets leverage simulated sensors (pinhole and stereo) to generate RGB and depth data, and we outline specific strategies used to handle scene diversity, pose recording, and structural alignment.
Gibson: This dataset covers 129 floors across 85 scenes. To ensure diversity, we randomize seed numbers during data generation, allowing for efficient image selection that maximizes spatial coverage with minimal redundancy. The image selection process is further detailed in Sec. B.1. We record precise camera poses using Habitat-sim and organize images into structured scenarios for downstream use.
HM3D: This dataset contains 755 annotated scenes. Since floor heights are not provided, we estimate them by clustering the Y-axis values of camera poses. This enables us to establish clear floor boundaries, which are essential for scene construction and co-visibility graph generation.
B Annotation generation details
B.1 Co-Visibility annotation overview
As described above, with the 3D assets and depth information provided in the dataset, the ground truth annotation of co-visibility between two images should theoretically be straightforward. However, this process is more complex than it seems. We follow the dataset generation process outlined in Sec. 3.3. First, while placing cameras and selecting images, the image set must remain sparse, maintaining both co-visibility space and maximum scene coverage. Second, as a reasoning task, the annotations should align with human perception while being automatically applicable to large-scale datasets. To ensure annotation accuracy and consistency, we curate a smaller, human-annotated dataset, which serves as a human reasoning baseline for benchmarking Co-VisiON and validating the correctness of the automatically generated dataset.
B.1.1 Automatic co-visibility annotation
Camera pose placing strategy. While we could place cameras anywhere within the navigable areas of a scene, we select positions near walls or furniture to better mimic human photography behavior. This strategy aims to emulate real-world scenarios, where photographers often position themselves near peripheral areas to maximize coverage. As shown in  Fig. I, the red-bordered regions between black obstacles and white navigable spaces indicate our preferred camera locations. The cameras are generally oriented away from walls to ensure comprehensive coverage, capturing a wide range of visual features that aid both neural network analysis and human interpretation.
Image selection criterion. First, to assess the scene coverage and the overlapping regions between any two images, we convert the depth image into a point cloud in global coordinates and use it in the subsequent scoring process. We select images in progressive iterations by adopting a scoring method to ensure we cover the entire scene with the fewest possible photos. In each iteration, $n_{c}$ randomly sampled candidates are generated based on the camera pose placing strategy where every candidate, $n_{c}^{i}$, possesses observations (RGB and Depth) along with the pose information of the agent and its sensors. We then choose the highest-scoring candidate based on the scoring function:
$S=\alpha\cdot O_{u}+\beta\cdot O_{p},$ (5)
where $O_{u}$ represents the newly explored region covered by the projected point cloud, while $O_{p}$ denotes the previously explored region recorded from previous iterations. The $\alpha$ and $\beta$ are set to 0.9 and 0.1, respectively. This configuration reflects a preference for camera poses that explore more of the uncovered area.
After selecting the best candidate from $n_{c}$ possible candidates based on the weighted score as in (5), we then remove the positions near the best candidate (or within some radius $r$) from the viable candidate selection area using equation (6):
$d(p_{i},p_{s})<=r$ (6)
[…]
B.1.2 Human co-visibility annotation
Besides developing the automatic annotation method, we also have human annotators manually label a subset of scenes. The human annotation mirrors how humans reason about spatial relationships and, therefore, serves two purposes: it helps assess the quality of automatic annotation, and more importantly, provides a human reasoning baseline for benchmarking Co-VisiON.
The process involves 6 scenes arbitrarily chosen from the automatically generated dataset from  Sec. B.1.1. To facilitate human annotation, we develop a website with a GUI that loads pairs of images within the same scene for the trained annotators to determine if they can reason the co-visibility of any pair of images, and click to label them accordingly. Images are uniformly sampled from the same scenario to form a pair and presented to the annotators.
Given images $I_{a}$, $I_{b}$ $\in\mathcal{I}$, the following criteria were considered for each pair:
  • Shared objects: If $I_{a}$ and $I_{b}$ share the view of the same objects such that an annotator can infer the two images’ relative pose, they are labeled as connected.
  • Object Continuity:
    If $I_{a}$ shows the left side of a sofa and $I_{b}$ shows the right side, even if the sofa is not fully visible in either image, the partial views align in a way that the viewer can perceive them as continuous, and spatially reason the images as parts of the same object.
  • Sub-Scene Relationship: one image may be a more zoomed-in portion of the other. Or an image may have details obstructed by objects in its view.
  • Featureless Surface: If the overlapping region of $I_{a}$ and $I_{b}$ is devoid of distinguishable features—such as a plain wall—the pair is labelled as not connected. The absence of features hampers the ability to establish a clear co-visibility relationship between the images.
Along with the above criterion, this manual annotation process undergoes a stringent cross-validation phase where each scene is annotated at least twice by different annotators. When it is challenging to identify relationships in ambiguous scenarios, the images may be marked for further review. The resulting sets of annotations are then compared and discussed until a complete agreement is reached among the annotators. In this way, we ensure the annotations of spatial relationships are as precise and robust as possible.
We will release both the automatically labeled and the manually labeled datasets to facilitate future research. The human annotation website will also be released so that anyone can extend the co-visibility reasoning task to a larger scale and for more diverse scenarios.
C Visualization of multiple scenarios from a single scene
We can simulate different scenarios within the same scene by manipulating the seed, resulting in distinct image combinations. In this supplementary, we present five scenarios generated for Scene Goff through this approach shown in Fig. II.
D Evaluation metrics
As discussed in Sec. 3.2, we use IoU (Intersection over Union) and AUC (Area Under Curve) as the evaluation criteria which are mathematically defined as following.
$\text{IOU}(\mathbf{A},\hat{\mathbf{A}})=\frac{\sum_{i}\sum_{j}\mathbf{A}_{ij}\land\hat{\mathbf{A}}_{ij}}{\sum_{i}\sum_{j}(\mathbf{A}_{ij}\lor\hat{\mathbf{A}}_{ij})+\epsilon}$ (7)
where $\mathbf{A}_{ij}$ and $\hat{\mathbf{A}}_{ij}$ stand for the binary elements at the $i$th row and $j$th column of the corresponding matrices of the ground truth and predicted co-visibility graphs, $\mathcal{G}$ and $\hat{\mathcal{G}}$. The $\land$ and $\lor$ represent element-wise AND and OR operations respectively. A small constant $\epsilon$ is added to the denominator to avoid zero division.
Another metric that we use is Area Under Curve (AUC). Formally, suppose $\hat{\mathbf{A}}_{\tau}$ is the binary predicted matrix given the threshold $\tau$, and $\tau$ is sampled from a discrete and ordered set $\{t_{1},t_{2},\ldots,t_{n}\}$ from range $[0,1]$. The AUC is computed as:
$\text{AUC}=\sum_{i=1}^{n-1}\frac{(\text{IOU}(\mathbf{A},\hat{\mathbf{A}}_{t_{i+1}})+\text{IOU}(\mathbf{A},\hat{\mathbf{A}}_{t_{i}}))}{2}\cdot(t_{i+1}-t_{i})$ (8)
E Details and graph definition in DUSt3R experiment
To evaluate the performance of co-visibility prediction under different graph structures, we experiment with several types of graphs in the DUSt3R benchmark. Each graph is defined as follows:
  • Complete graph refers to a graph where every pair of images is directly connected. However, due to its resource-intensive nature, it becomes less practical for large-scale reconstructions.
  • Co-visibility graph is resource-efficient and maintains geometric coherence, making it a viable alternative to the Complete strategy.
  • Star graph consists of a central frame connecting all other images. This approach emphasizes a central perspective and proves valuable when a single viewpoint dominates the scene.
  • Ground Truth graph connects a pair of images in its graph if they are within a predefined proximity. This graph relies on the absolute pose obtained from the ground truth for each pair of images.
In addition, we present a comparison of these graphs’ performance in terms of 3D reconstruction accuracy and computational cost. As shown in Fig. III, we visualize how each graph affects image pairing during reconstruction. The advantages and limitations of each graph type, including aspects such as memory usage and scalability, are discussed in detail in Sec. 5.1. This analysis helps justify our choice of the co-visibility graph as the optimal solution for balancing accuracy and computational efficiency.
(a) Star - Scene 1
(b) Covision - Scene 1
(c) Complete - Scene 1
(d) GT - Scene 1
(e) Star - Scene 2
(f) Covision - Scene 2
(g) Complete - Scene 2
(h) GT - Scene 2
Figure III: Two examples of 3D reconstruction results by DUSt3R, each based on distinct graphs.
F Detailed definition of different graphs used in CroCo experiment
We train a model to evaluate its performance over the co-visible regions defined by different graph structures, including: (1) High Co-visibility Graph, (2) Co-visibility Graph, and (3) Random Graph.
  • High co-visibility graph: This graph is created using the pairs of images from a scenario with high spatial and visual overlap ($\geq$50%, or IoU i). Using a threshold i, it selects pairs with significant overlap, focusing on high-confidence spatial relationships. This graph aims to test the CroCo model’s capability to reconstruct a masked target image using a reference image from a different perspective, particularly emphasizing close spatial relationships.
  • Co-visibility graph: The Co-Visibility graph of a scenario is the same graph as discussed in Sec. 3.1 and Sec. 3.3. This usually contains image pairs with both high and low visual/spatial overlap connected by an edge. From these pairs, we randomly select the same number of edges as in the High Co-visibility Graph. This selection ensures a balanced comparison, containing a mixture of high and low co-visibility pairs. This graph is useful in testing the CroCo model’s ability to reconstruct images from a variety of perspectives, leveraging both high and low co-visibility information.
  • Random graph: The Random graph is constructed by selecting all possible image pairs within a scenario, including those with little to no spatial or visual overlap. A subset of these pairs is randomly selected, ensuring the number of edges matches that of the High Co-visibility Graph. This diverse selection, which includes pairs with minimal or no overlap, tests the CroCo model’s ability to handle a variety of image combinations, challenging it to perform well across more unpredictable scenarios.
G Ablation Study of CoVis
G.1 Effect of noisy mask on Covis
Mask prediction plays a crucial role in CoVis, as demonstrated in  Tab. 2. In this subsection, we explore how the accuracy of ground truth co-visible masks—ranging from accurate to noisy—affects the model’s performance on the test set.
To simulate noisy supervision, we introduce boundary-localized Gaussian perturbations to the ground truth binary masks. First, a Laplacian filter is used to detect boundaries between foreground and background. Gaussian noise is then added to these boundary pixels, with standard deviations of $\sigma=0.01$ for low noise and $\sigma=0.5$ for high noise. The perturbed mask is re-binarized using a threshold of $\tau=0.5$, while retaining all original positive pixels to avoid removing valid co-visible regions. If the original mask is entirely empty, no noise is applied.
When provided with low levels of noise in the co-visibility masks, the performance on both IoU and AUC remains comparable or even slightly improves due to the regularizing effect of noise, akin to data augmentation. However, as shown in Tab. I, excessive noise leads to a significant drop in performance, highlighting the importance of maintaining reasonable mask quality.
Table I: Comparison of different mask types with varying noise levels in Covis’ test performance (IoU means Graph IoU).
Mask Type Noise Gibson HM3D
IoU AUC IoU AUC
No Mask - 0.52 0.49 0.48 0.42
Supervised Mask None 0.59 0.57 0.57 0.56
Supervised Mask Low 0.59 0.56 0.59 0.55
Supervised Mask High 0.56 0.50 0.50 0.48
Applied GT Mask - 0.74 0.72 0.71 0.69
G.2 Zero-shot Evaluation
We further evaluate the generalization ability of our Covis by training on one dataset (either Gibson or HM3D) and testing on both Gibson and HM3D individually. As shown in Tab. II, Multi-view Covis consistently outperforms pairwise Covis in both in-domain and zero-shot settings, demonstrating better robustness to dataset’s domain shifts. We also evaluate the predicted masks using the standard IoU metric (not graph IoU), and achieve an average IoU of 67.3% against the binarized ground-truth masks.
Table II: Zero-shot evaluation of multi-view Covis and pairwise Covis across different training and test sets.
Train Test Multi-view Covis Pairwise Covis
IoU AUC IoU AUC
Gibson Gibson 0.56 0.52 0.51 0.47
Gibson HM3D 0.51 0.48 0.48 0.44
HM3D Gibson 0.58 0.53 0.55 0.50
HM3D HM3D 0.52 0.48 0.50 0.46
G.3 Qualitative Result of Mask Prediction
We also visualize the ground-truth co-visibility masks between image pairs; which serve as additional supervision signals during training, and the binary co-visibility masks learned by Covis model in Fig. IV.
(a) Scene 1 with Green Highlighted Overlap from Scene 2
(b) Scene 2 with Green Highlighted Overlap from Scene 1
(c) Binarized GT mask for (a)
(d) Binary predicted mask for (a)
Figure IV: Visualization of the co-visible region between a pair of images (a) and (b), with (c) showing the binarized ground truth mask and (d) the predicted mask on the image from (a).
H Failure mode analysis
We add more baselines: Gemini-2.0-Flash [57] (a recent VLM with better spatial understanding), SpatialRGPT [11] (a recent Spatial-VLM), Covis-freeze (a variant with the CrocoV2 backbone frozen except for the last two layers) and SigLIP (a contrastive learning baseline). All baselines were evaluated on 100 Gibson test scenarios (5 per scene), except for Gemini-2.0-Flash, which was run on 20 scenarios using the first scenario from each scene due to limited time and compute during rebuttal. A full-scale benchmark for Gemini-2.0-Flash will be made available upon paper acceptance. Although we cannot and were not meant to be exhaustive, we added the above mentioned more recent baselines in our camera-ready version because of their better spatial reasoning capabilities. For the GPT-4o and Gemini 2.0 Flash baselines, we provide image pairs along with a prompt that explicitly asks the model to reason about co-visibility between the two images. SpatialRGPT, is built on Vila[31], which supports multi-image input. We feed image pairs into its visual encoder and use the fine-tuned LLaMA3-70B to reason about co-visibility from the encoded tokens. We use the same prompt for GPT-4o, Gemini 2.0 Flash, and SpatialRGPT, as shown inFig. V. For the failure mode analysis, we categorize test scenes into easy, medium, and hard levels based on scene sparsity and pairwise overlap (see the caption of Tab. III).
Figure V: An example of the prompt used in our experiments. A unified prompt is adopted across all vision-language models, including GPT-4o, Gemini-2.0-Flash, and SpatialRGPT.
Table III: Best IoU performance across different difficulty levels under two evaluation protocols: (1) Image Overlap, categorized based on the overlap ratio of each image pair — easy: $\geq 50\%$, medium: $\geq 10\%$ and $<50\%$, hard: $<10\%$. (2) Scene Sparsity, defined by the average pairwise overlap within each scene — easy: $\geq 10\%$, medium: $\geq 4\%$ and $<10\%$, hard: $<4\%$. Top-3 performers in each column are highlighted using the color scheme consistent with Table 2 in the main paper.
Setting Method Easy Med. Hard Avg.
Image Overlap (edge level) GPT-4o 0.97 0.92 0.34 0.63
Gemini-2.0-Flash 1.00 0.99 0.14 0.42
Qwen2.5-VL 72b 1.00 0.99 0.14 0.41
SigLIP 1.00 1.00 0.13 0.35
SpatialRGPT 1.00 0.99 0.13 0.35
Covis 0.99 0.88 0.24 0.59
Covis (freeze) 1.00 0.89 0.30 0.61
Scene Sparsity (graph level) GPT-4o 0.83 0.64 0.57 0.63
Gemini-2.0-Flash 0.75 0.39 0.28 0.42
Qwen2.5-VL 72b 0.77 0.40 0.28 0.41
SigLIP 0.73 0.38 0.26 0.35
SpatialRGPT 0.72 0.38 0.26 0.35
Covis 0.80 0.63 0.52 0.59
Covis (freeze) 0.81 0.65 0.54 0.61
Discussion. In Tab. III, GPT-4o outperforms other VLMs. Thanks to R3’s suggestion, Covis (freeze) slightly improves over the original. Both Covis and GPT-4o achieve the most significant improvements over all baselines in hard case scenarios. These hard cases highlight that limited viewpoint overlap and scene sparsity are the primary sources of difficulty.
I Sim2Real
In this experiment, our aim is to demonstrate that the Co-VisiON dataset exhibits a small domain gap compared to real-world environments, such as the AVD dataset [3]. We have tested Covis, ViT, VGG, Resnet, Contrastive, and NetVlad methods on the AVD dataset using pretrained model from Co-VisiON dataset. We observe AUC metric results in Tab. IV to be comparable to those in Tab. 2. Similarly, from Fig. VI, we can see that the co-visibility graph predicted using the Covis closely resembles the manually labelled topology graph.
(a) Manually Labeled Topology
(b) Covis Predicted Topology
Figure VI: Comparison of manually labeled and Covis predicted Topologies. For clarity, we show 10 sampled images to provide an illustrative example.
Table IV: Zero-shot comparison of AUC values (%) for baseline models pretrained on the Co-VisiON dataset and evaluated on real-world data.
Baseline Covis ViT VGG ResNet Contrastive NetVlad
AUC 0.61 0.52 0.33 0.32 0.31 0.22